OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL AND FAN Below are the steps to solve an equation: step 1: [x - 5] + 2 = 5 step 2: [x - 5] = 5 - 2 step 3: [x - 5] = 3 which of the following is a correct next step to solve the equation? A. x + 5 = -3 B. -x - 5 = 3 C. x - 5 = -3 D. x + 5 = 3

OpenStudy (hesan):

i think all the given options are wrong.

OpenStudy (here_to_help15):

\[\left| x-5 \right| +2=5\rightarrow \left| x-5 \right|=5-2\rightarrow \left| x-5 \right|=3\]

OpenStudy (texaschic101):

| x - 5 | = 3 x - 5 = 3 and -(x - 5) = 3 which is the same as -x + 5 = 3....oh..thats not an answer choice -(x - 5) = 3 -x + 5 = 3 --- multiply by -1 to make x positive x - 5 = -3....I am going with C